Arrange the following oxoacids in order of decreasing acid strength.

Part A Arrange the following oxoacids in order of decreasing acid strength.
Rank from strongest to weakest acid. To rank items as equivalent, overlap them.
1. HClO2
2. HClO3
3. HBrO
4. HClO

Part B Arrange the following carboxylic acids in order of decreasing acid strength.
1. CHCl2COOH
2. CH2ClCOOH
3. CH3COOH
4. CH3CH2COOH

Part C Arrange the following amines in order of decreasing base strength.
1. CH3NH2
2. NH3
3. NH2Br
4. (CH3)2NH
Answer #1

General guidance

Concepts and reason

Any substance can be classified into an acid and a base in chemistry. An acid is a substance which donates a hydrogen ion (H+)\\left( {{{\\rm{H}}^{\\rm{ + }}}} \\right) or proton in solutions and a base is a substance which donates a hydroxide ion (OH\u2212)\\left( {{\\rm{O}}{{\\rm{H}}^{\\rm{ - }}}} \\right)in solution. Furthermore, these can be classified by using Arrhenius theory, Bronsted and Lowry theory and Lewis acid base theory.

Fundamentals

Classification of substances into an acid and a base is based on the following criteria

\uf0d8Oxidation number of the central atom should be noted in oxo acids since the acidity increases with increase in oxidation number of the central atom.

\uf0d8The presence of more electronegative atoms increases the acidity of the substance.

\uf0d8Presence of more electron donating groups increases the basicity of chemical substances

First Step | All Steps | Answer Only

Step-by-step

Step 1 of 7

The given acids are as follows.

HClO3,HClO2,HClO,HBrO{\\rm{HCl}}{{\\rm{O}}_{\\rm{3}}}{\\rm{,HCl}}{{\\rm{O}}_{\\rm{2}}}{\\rm{,HClO,HBrO}}

Explanation | Common mistakes | Hint for next step

The central atoms of the given acids are Cl{\\rm{Cl}} and Br{\\rm{Br}} since they are the most electronegative atoms in the given acids. The structure of the given acids is given below:

Step 2 of 7

The given acids are as follows.

le:31-7011-C

Explanation | Common mistakes | Hint for next step

The oxidation number of the central atom is calculated for all the given acids by considering the oxidation number of oxygen atom is -2 and hydroxide (OH{\\rm{OH}}) is -1.The oxidation state for acids are calculated as shown below:

HClO3x+(\u22126+1)=0(wherexdenotestheoxidationstateofCl)x\u22125=0x=5HencetheoxidationstateofClis+5\\begin{array}{l}\\\\{\\rm{HCl}}{{\\rm{O}}_{\\rm{3}}}\\\\\\\\{\\rm{x}}\\,{\\rm{ + }}\\,\\left( { - {\\rm{6 + 1}}} \\right)\\,\\,{\\rm{ = }}\\,\\,{\\rm{0}}\\,\\,\\left( {{\\rm{where}}\\,{\\rm{x}}\\,{\\rm{denotes}}\\,{\\rm{the}}\\,{\\rm{oxidation}}\\,{\\rm{state}}\\,{\\rm{of}}\\,{\\rm{Cl}}} \\right)\\\\\\\\{\\rm{x}}\\,\\, - {\\rm{5}}\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,{\\rm{ = }}\\,\\,{\\rm{0}}\\\\\\\\\\,\\,\\,\\,\\,\\,\\,{\\rm{x}}\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,\\,{\\rm{ = }}\\,{\\rm{5}}\\\\\\\\{\\rm{Hence}}\\,{\\rm{the}}\\,{\\rm{oxidation}}\\,{\\rm{state}}\\,{\\rm{of}}\\,{\\rm{Cl}}\\,{\\rm{is}}\\,{\\rm{ + 5}}\\\\\\end{array}